If the price of rice per Kg increases from Rs 200 to Rs 300, the quantity demand reduced from 10 Kg units to 8 Kg. Keeping the answer in view if price change by 7%, what happen to Quantity demand of the commodity
1
Expert's answer
2020-12-15T07:01:39-0500

Solution:

Calculate the price elasticity of demand (Ed):


"E_{d} =\frac{\%\;change\; in \;quantity \;demand}{\%\;change\; in \;price}"


First, calculate the % change in quantity demand (Qd):

"Q_{d} =\frac{8-10}{(8+10)\div 2 } =\frac{-2}{9} =-0.22"


Calculate the % change in price:

"\%\;change\; in \;price =\frac{300-200}{(300+200)\div 2 } =\frac{100}{250} =0.40"


"Price \; elasticity\; of \; demand \; (E_{d}) =\frac{-0.22}{0.40} =-0.5556"


If the price changes by 7%:


"E_{d} =\frac{\%\;change\; in \;quantity \;demand}{\%\;change\; in \;price}"


"-0.5556 =\frac{\%\;change\; in \;quantity \;demand}{7}"


"\%\;change\; in \;quantity \;demand =-0.5556\times 7"

"=-3.89\%"


The Quantity demand of the commodity will fall by 3.89% when the price changes by 7%